# Computing algebraic Belyi functions on Bring’s curve

### v2

@@ Abstract.
In this paper, we explicitly compute two kinds of algebraic Belyi functions on Bring’s curve. One is related to a congruence subgroup of ${\rm SL}_{2}(\mathbb{Z})$ and the other is relate to a congruence subgroup of the triangle group $\Delta(2,4,5)\subset{\rm SL}_{2}(\mathbb{R})$. To carry out the computation, we use elliptic cusp forms of weight 2 for the former case while the automorphism group of Bring’s curve for the latter case. We also discuss a suitable base field (a number field) for isomorphisms between Hulek-Craig’s curve, Bring’s curve, and its another algebraic model obtained as a modular curve.
@@ 1. Introduction
The main result of Belyi [1] shows that a complex projective smooth curve $C$ has a model, as an algebraic curve, defined over a number field if and only if there exists a non-constant function $\beta:C\longrightarrow\mathbb{P}^{1}$ ramified at most over $\{0,1,\infty\}$. In this paper, we call such a function $\beta$ an algebraic Belyi function and more precisely, an algebraic Belyi function defined over a number field $K$ if it is defined over $K$. Thus, any projective smooth geometrically connected curve over a number field is endowed with an algebraic Belyi function. However, finding such a $\beta$ explicitly for a given $C$ is another matter. In fact, many people have studied explicit computation of algebraic Belyi functions (cf. [23],[21] among others). Belyi’s result can also be understood by hyperbolic uniformization and there are two types of such uniformization:
@@
    (Non-compact case) There exist a finite index subgroup $\Gamma$ of ${\rm SL}_{2}(\mathbb{Z})$ and a biholomorphism $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
    $X(\Gamma)\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Gamma\tau\mapsto{\rm SL}_{2}(\mathbb{Z})\tau\mapsto j(\tau)$
    where $j$ is Klein’s $j$-invariant function so that $j(\tau)=q^{-1}+744+196884q+\cdots,\ q:=e^{2\pi\sqrt{-1}\tau},\ \tau\in\mathbb{H}:=\{\tau\in\mathbb{C}\ |\ {\rm Im}(\tau)>0\}$.
@@
    (Compact case) There exist a finite index subgroup $\Delta$ of some trignale group $\Delta(a,b,c)\subset{\rm SL}_{2}(\mathbb{R})$ and a biholomorphism $X_{\Delta}:=\Delta\backslash\mathbb{H}\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
@@
We refer [17, p.71, Theorem] for the non-compact case and [24, Section 3.3, Theorem 3] for the compact case.
In this paper, using the above two interpretation, we will compute two algebraic Belyi functions on Bring’s curve defined by
@@
inside $\mathbb{P}^{4}$ with homogeneous coordinates $[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]$. The readers may consult with Theorem 2.2 and Theorem 3.1 as the main results of this paper.
Let us explain about some key points for the main results. Let $S_{5}$ be the symmetric group on 5 letters. It is known that
@@
    (non-compact case [9, p.82-83]) $B(\mathbb{C})\simeq X(\Gamma_{B})$ where $\Gamma_{B}\subset{\rm SL}_{2}(\mathbb{Z})$ is explicitly given in the next section (see also Section 4), and
@@
For the non-compact case, we compute the $j$-function in terms of the modular functions on $X(\Gamma_{B})$ by regarding $j$ and modular functions with algebraic coordinates.
For the compact case, we first use the uniformization to conclude the natural quotient map $B(\mathbb{C})\mapsto B(\mathbb{C})/{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))$ is an algebraic Belyi function. As a projective smooth complex curve, Bring’s curve is a unique curve of genus 4 with $S_{5}$ as the automorphism group. On the other hands, $S_{5}$ acts properly on the coordinates of $B$ and if we view $B$ as an algebraic curve over $\mathbb{Q}$, we have ${\rm Aut}_{\mathbb{Q}}(B)\supset S_{5}$ where ${\rm Aut}_{\mathbb{Q}}$ stands for the group of all automorphims defined over $\mathbb{Q}$. However, $S_{5}\subset{\rm Aut}_{\mathbb{Q}}(B)\subset{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))\simeq S_{5}$. Thus, ${\rm Aut}_{\mathbb{Q}}(B)=S_{5}$. Hence, the natural quotient map decent to the quotient map $B\longrightarrow B/{\rm Aut}_{\mathbb{Q}}(B)\simeq\mathbb{P}^{1}_{\mathbb{Q}}$ which is an algebraic Belyi function defined over $\mathbb{Q}$.
We will organize this paper as follows. In Section 2, we study another model of $B$ which is explicitly given as an algebraic model of the modular curve $X(\Gamma_{B})$. By using elliptic cusp forms of weight 2 and theory of canonical curves, we compute an explicit defining equation over $\mathbb{Q}$ for $X(\Gamma_{B})$. In Section 3, we will compute the quotient map $B\mapsto B/{\rm Aut}_{\mathbb{Q}}(B)$ explicitly. Finally, in Section 4, we relate our algebraic model of $X(\Gamma_{B})$ with Hulek-Craig’s curve and discuss a number field $K$ such that these curves are isomorphic over $K$ to Bring’s curve. This shows that Bring’s curve inherits two interesting algebraic Belyi functions.
Acknowledgments. We would like to thank Professors K. Hulek and H. Braden for helpful discussions.
@@ 2. Bring’s curve as a modular curve
In this section, we compute the Belyi function on the Bring’s curve $B$ as the modular curve $X(\Gamma_{B})$ introduced as below. It will be explained in Section 4 that $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})$. Thus, we may work on $X(\Gamma_{B})$ to compute the Belyi function.
We refer [5] for the basic facts of elliptic modular forms and modular curves. For each positive integer $N$, we denote by $\Gamma(N)\supset\Gamma_{0}(N)\supset\Gamma_{1}(N)$ the three kinds of congruence subgroups inside ${\rm SL}_{2}(\mathbb{Z})$ introduced in [5, p.13, Definition 1.2.1]. For such a congruence subgroup $\Gamma$, we define the (compact) modular curve $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))$ and put $X_{0}(N)=X(\Gamma_{0}(N))$ for simplicity. We denote by $S_{2}(\Gamma)$ the space of all cusp forms of weight 2 with respect to $\Gamma$. It is known that we have an isomorphism $S_{2}(\Gamma)\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{0}(X(\Gamma),\Omega^{1}_{X(\Gamma)}),\ f(\tau)\mapsto f(\tau)d\tau$ where the right hand side stands for the space of all holomorphic 1-forms on $X(\Gamma)$.
@@ 2.1. Defining equations over $\mathbb{Q}$ of $X(\Gamma_{B})$
It is well-known (cf. [6, Proposition 2.1]) that $\beta_{\Gamma_{B}}$ is a Belyi function which is ramified only at $0,1,\infty$. Since each of $X(\Gamma_{B})$ and $X({\rm SL}_{2}(\mathbb{Z}))$ has a structure as a projective smooth algebraic curve defined over $\mathbb{Q}$, the Belyi function $\beta_{\Gamma_{B}}$ can be seen as an algebraic function $\beta^{{\rm alg}}_{\Gamma_{B}}$ over $\mathbb{Q}$. Henceforth, we will explicitly compute $\beta^{{\rm alg}}_{\Gamma_{B}}$.
@@
(2.2) $X:\left\{\begin{array}[]{l}-(x^{2}+y^{2})+2(z^{2}+w^{2})=0\\ x^{3}+xy^{2}+2y^{3}-4xz^{2}-8xzw=0\end{array}\right.$
@@
(2.3) $\begin{array}[]{l}f_{1}(\tau):=f_{{\rm 50.2.a.a}}(q)=q-q^{2}+q^{3}+q^{4}-q^{6}+2*q^{7}-\cdots\\ f_{2}(\tau):=f_{{\rm 50.2.a.b}}(q)=q+q^{2}-q^{3}+q^{4}-q^{6}-2*q^{7}+q^{8}-\cdots\\ f_{3}(\tau):=\frac{1}{2}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})+f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q-q^{4}-q^{6}+\cdots\\ f_{4}(\tau):=\frac{1}{2\sqrt{-1}}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})-f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q^{2}+q^{3}-2q^{7}-q^{8}-\cdots\end{array}$
where each label $\ast$ of $f_{\ast}$ corresponds to the corresponding datum on [13] and we can read off the Fourier coefficients up to $q^{99}$ for which it is enough for the computation below. Since Bring’s curve is non-hyperelliptic and of genus 4, its canonical divisor is very ample. Hence, by noting $S_{2}(\Gamma_{B})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{0}(X(\Gamma_{B}),\Omega^{1}_{X(\Gamma_{B})}),\ f(\tau)\mapsto f(\tau)d\tau$, we have the canonical embedding:
@@
By using Petri’s theorem as exlpained in [20, Section 2.2] for $X_{0}(N)$ (which also applies to $\Gamma_{B}$), the image of $\Phi$ is given by the intersection of a quadratic equation and a cubic equation in $\mathbb{P}^{3}$. The claim follows by determining the coefficients of those equations by using (2.3) (see [20] for details in the computation). ∎
2.2. An algebraic Belyi function on $X$
In this subsection, we compute the algebraic function over $\mathbb{Q}$ corresponding to the quotient map $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle{\rm SL}_{2}(\mathbb{Z})\tau\mapsto j(\tau)}}{{\simeq}}\mathbb{P}^{1}(\mathbb{C})$ by using the intermediate modular curve $X_{0}(50)$ so that the quotient map factors through it:

### v3

@@ Abstract.
In this paper, we explicitly compute two kinds of algebraic Belyi functions on Bring’s curve. One is related to a congruence subgroup of ${\rm SL}_{2}(\mathbb{Z})$ and the other is related to a congruence subgroup of the triangle group $\Delta(2,4,5)\subset{\rm SL}_{2}(\mathbb{R})$. To carry out the computation, we use elliptic cusp forms of weight 2 for the former case and the automorphism group of Bring’s curve for the latter case. We also discuss a suitable base field (a number field) for describing isomorphisms between Hulek-Craig’s curve, Bring’s curve, and another algebraic model obtained as a modular curve.
@@ 1. Introduction
The main result of Belyi [1] shows that a complex projective smooth curve $C$ has a model, as an algebraic curve, defined over a number field if and only if there exists a non-constant function $\beta:C\longrightarrow\mathbb{P}^{1}$ ramified at most over $\{0,1,\infty\}$. In this paper, we call such a function $\beta$ an algebraic Belyi function and more precisely, an algebraic Belyi function defined over a number field $K$ if it is defined over $K$. Thus, any projective smooth geometrically connected curve over a number field is endowed with an algebraic Belyi function. However, finding such a $\beta$ explicitly for a given $C$ is another matter. In fact, the explicit computation of algebraic Belyi functions has been studied extensively (cf. [23],[21] among others). Belyi’s result can also be understood via hyperbolic uniformization and there are two types of uniformization:
@@
    (Non-compact case) There exists a finite index subgroup $\Gamma$ of ${\rm SL}_{2}(\mathbb{Z})$ and a biholomorphism $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
    $X(\Gamma)\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Gamma\tau\mapsto{\rm SL}_{2}(\mathbb{Z})\tau\mapsto\frac{1}{1728}j(\tau)$
    where $j$ is Klein’s $j$-invariant function, with expansion $j(\tau)=q^{-1}+744+196884q+\cdots,\ q:=e^{2\pi\sqrt{-1}\tau},\ \tau\in\mathbb{H}:=\{\tau\in\mathbb{C}\ |\ {\rm Im}(\tau)>0\}$.
@@
    (Compact case) There exists a finite index subgroup $\Delta$ of some triangle group $\Delta(a,b,c)\subset{\rm SL}_{2}(\mathbb{R})$ and a biholomorphism $X_{\Delta}:=\Delta\backslash\mathbb{H}\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
@@
We refer to [17, p.71, Theorem] for the non-compact case and [24, Section 3.3, Theorem 3] for the compact case.
In this paper, using the above two interpretations, we compute two algebraic Belyi functions on Bring’s curve defined by
@@
inside $\mathbb{P}^{4}$ with homogeneous coordinates $[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]$. The symmetric group $S_{5}$ of degree 5 acts on $B$ by permutation of coordinates and it is known that $S_{5}$ is the full automorphism group of $B$. We refer to [25, Example 5.5 and Proposition 5.6] for another aspect of Bring’s curve.
A key ingredient in computing the Belyi maps is to relate Bring’s curve with a certain modular curve and Hulek-Craig’s curve. We briefly introduce these curves and the details are given in the corresponding sections. Let $X(\Gamma_{B})^{{\rm alg}}$ be the algebraic model over $\mathbb{Q}$ of the compactified modular curve $X(\Gamma_{B}):=\overline{\Gamma_{B}\backslash\mathbb{H}}$ which is given by
(1.2) $X(\Gamma_{B})^{{\rm alg}}:\left\{\begin{array}[]{l}-(x^{2}+y^{2})+2(z^{2}+w^{2})=0\\ x^{3}+xy^{2}+2y^{3}-4xz^{2}-8xzw=0\end{array}\right.$
inside $\mathbb{P}^{4}$ (see Section 2). Here $\Gamma_{B}:=\Gamma_{0}(2)\cap\Gamma(5)$ is a congruence subgroup in ${\rm SL}_{2}(\mathbb{Z})$. Let $\beta^{{\rm alg}}_{\Gamma_{B}}:X(\Gamma_{B})^{{\rm alg}}\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}}$ be an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C}),\tau\mapsto\frac{1}{1728}j$ (see Theorem 2.2).
On the other hand, Hulek-Craig’s curve $X^{{\rm HC}}$ over $\mathbb{Q}$ is defined by the desingularization of a projective geometrically irreducible singular curve
inside $\mathbb{P}^{2}$ with the coordinates $[x_{1}:y_{1}:z_{1}]$. This curve has studied in [9] and [4]. Hulek-Craig’s curve $X^{{\rm HC}}$ appears in the context of the Horrocks-Mumford bundle [9]. Through its symmetries, the latter is also closely related to the group $\Gamma(5)$.
Then we prove the following result:
Theorem 1.1.
(Theorem 4.1) Keep the notation as above. Let $\zeta_{5}$ be a primitive fifth root of unity.
@@
    Bring’s curve $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to Hulek-Craig’s curve $X^{{\rm HC}}$;
  2. (2)
    Hulek-Craig’s curve $X^{{\rm HC}}$ is isomorphic over $\mathbb{Q}$ to $X(\Gamma_{B})^{{\rm alg}}$.
We note that the first part of Theorem 1.1 is due to [2, Proposition 2.9]. The main results of this paper are Theorem 2.2 and Theorem 3.1 and they are summarized as follows:
Theorem 1.2.
Bring curve has two algebraic Belyi functions which are given by the following contexts:
  1. (1)
    (non-compact type) The composition of $B\stackrel{{\scriptstyle\mathbb{Q}(\zeta_{5})}}{{\simeq}}X^{{\rm HC}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\simeq}}X(\Gamma_{B})^{{\rm alg}}$ and $\beta^{{\rm alg}}_{\Gamma_{B}}:X(\Gamma_{B})^{{\rm alg}}\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}}$ gives an algebraic Belyi map over $\mathbb{Q}(\zeta_{5})$ on $B$.
  2. (2)
    (compact type) The natural quotient $B\longrightarrow B/S_{5}$ gives an algebraic Belyi map over $\mathbb{Q}$ on $B$. It is an algebraic model over $\mathbb{Q}$ of the natural quotient map $X_{\Delta}\longrightarrow X_{\Delta(3,4,5)}\simeq\mathbb{P}^{1}(\mathbb{C})$ induced by the inclusion of a certain subgroup $\Delta$ of $\Delta(3,4,5)$.
Let us explain about some key points for the main results. It is known that
  1. (1)
    (non-compact case [9, p.82-83]) $B(\mathbb{C})\simeq X(\Gamma_{B})$ as a Riemann surface, and
@@
For the non-compact case, we compute an algebraic model of the $j$-function in terms of the modular functions on $X(\Gamma_{B})$ by regarding $j$ and modular functions with algebraic coordinates.
For the compact case, we first use the uniformization to conclude that the natural quotient map $B(\mathbb{C})\mapsto B(\mathbb{C})/{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))$ is an algebraic Belyi function. As a projective smooth complex curve, Bring’s curve is the unique curve of genus 4 with $S_{5}$ as the automorphism group. On the other hands, $S_{5}$ acts properly on the coordinates of $B$ and if we view $B$ as an algebraic curve over $\mathbb{Q}$, we have ${\rm Aut}_{\mathbb{Q}}(B)\supset S_{5}$ where ${\rm Aut}_{\mathbb{Q}}$ stands for the group of all automorphims defined over $\mathbb{Q}$. However, $S_{5}\subset{\rm Aut}_{\mathbb{Q}}(B)\subset{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))\simeq S_{5}$. Thus, ${\rm Aut}_{\mathbb{Q}}(B)=S_{5}$. Hence, the natural quotient map decent to the quotient map $B\longrightarrow B/{\rm Aut}_{\mathbb{Q}}(B)\simeq\mathbb{P}^{1}_{\mathbb{Q}}$ which is an algebraic Belyi function defined over $\mathbb{Q}$.
We will organize this paper as follows. In Section 2, we study another model of $B$ which is explicitly given as an algebraic model of the modular curve $X(\Gamma_{B})$. By using elliptic cusp forms of weight 2 and the theory of canonical curves, we compute an explicit defining equation over $\mathbb{Q}$ for $X(\Gamma_{B})$. In Section 3, we will compute the quotient map $B\mapsto B/{\rm Aut}_{\mathbb{Q}}(B)$ explicitly. Finally, in Section 4, we relate our algebraic model of $X(\Gamma_{B})$ with Hulek-Craig’s curve and discuss a number field $K$ such that these curves are isomorphic over $K$ to Bring’s curve. This shows that Bring’s curve inherits two interesting algebraic Belyi functions.
Acknowledgments. We would like to thank Professors K. Hulek and H. Braden for helpful discussions. We also thank Natalia Amburg for kindly informing us an article [25]. Finally, we would like to give special thanks to the referee, whose suggestions have greatly improved the presentation and readability of this paper.
@@ 2. Bring’s curve as a modular curve
In this section, we compute the Belyi function on Bring’s curve $B$ as the modular curve $X(\Gamma_{B})$ defined below. It will be explained in Section 4 that $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})$. Thus, we may work on $X(\Gamma_{B})$ to compute the Belyi function.
We refer to [5] for the basic facts of elliptic modular forms and modular curves. For each positive integer $N$, we denote by $\Gamma(N)\supset\Gamma_{0}(N)\supset\Gamma_{1}(N)$ the three kinds of congruence subgroups inside ${\rm SL}_{2}(\mathbb{Z})$ introduced in [5, p.13, Definition 1.2.1]. For such a congruence subgroup $\Gamma$, we define the (compact) modular curve $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))$ and put $X_{0}(N)=X(\Gamma_{0}(N))$ for simplicity. We denote by $S_{2}(\Gamma)$ the space of all cusp forms of weight 2 with respect to $\Gamma$. It is known that we have an isomorphism $S_{2}(\Gamma)\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{0}(X(\Gamma),\Omega^{1}_{X(\Gamma)}),\ f(\tau)\mapsto f(\tau)d\tau$ where the right hand side stands for the space of all holomorphic 1-forms on $X(\Gamma)$.
@@ 2.1. Defining equations over $\mathbb{Q}$ of $X(\Gamma_{B})$
It is well-known (cf. [6, Proposition 2.1]) that $\beta_{\Gamma_{B}}$ is a Belyi function ramified only at $0,1,\infty$. Since each of $X(\Gamma_{B})$ and $X({\rm SL}_{2}(\mathbb{Z}))$ has a structure as a projective smooth algebraic curve defined over $\mathbb{Q}$, the Belyi function $\beta_{\Gamma_{B}}$ can be seen as an algebraic function $\beta^{{\rm alg}}_{\Gamma_{B}}$ over $\mathbb{Q}$. Henceforth, we will explicitly compute $\beta^{{\rm alg}}_{\Gamma_{B}}$.
@@
(2.2) $X(\Gamma_{B})^{{\rm alg}}:\left\{\begin{array}[]{l}-(x^{2}+y^{2})+2(z^{2}+w^{2})=0\\ x^{3}+xy^{2}+2y^{3}-4xz^{2}-8xzw=0\end{array}\right.$
@@
(2.3) $\begin{array}[]{l}f_{1}(\tau):=f_{{\rm 50.2.a.a}}(q)=q-q^{2}+q^{3}+q^{4}-q^{6}+2q^{7}-\cdots\\ f_{2}(\tau):=f_{{\rm 50.2.a.b}}(q)=q+q^{2}-q^{3}+q^{4}-q^{6}-2q^{7}+q^{8}-\cdots\\ f_{3}(\tau):=\frac{1}{2}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})+f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q-q^{4}-q^{6}+\cdots\\ f_{4}(\tau):=\frac{1}{2\sqrt{-1}}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})-f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q^{2}+q^{3}-2q^{7}-q^{8}-\cdots\end{array}$
where each label $\ast$ of $f_{\ast}$ corresponds to the corresponding datum on [13] and we can read off the Fourier coefficients up to $q^{99}$ for which it is enough for the computation below. Readers may visit the website https://www.lmfdb.org/ModularForm/GL2/Q/holomorphic/ linked in [13] and enter 50 for the level, and 2 for the weight, and either 1 or 2 for the character order. Then all labelled forms as above will show up individually.
Since Bring’s curve is non-hyperelliptic and of genus 4, its canonical divisor is very ample. Hence, by noting $S_{2}(\Gamma_{B})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{0}(X(\Gamma_{B}),\Omega^{1}_{X(\Gamma_{B})}),\ f(\tau)\mapsto f(\tau)d\tau$, we have the canonical embedding:
@@
By using Petri’s theorem as explained in [20, Section 2.2] for $X_{0}(N)$ (which also applies to $\Gamma_{B}$), the image of $\Phi$ is given by the intersection of a quadratic equation and a cubic equation in $\mathbb{P}^{3}$. The claim follows by determining the coefficients of those equations by using (2.3); see [20] for computational details. ∎
2.2. An algebraic Belyi function on $X(\Gamma_{B})^{{\rm alg}}$
In this subsection, we compute the algebraic function over $\mathbb{Q}$ corresponding to the quotient map $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle{\rm SL}_{2}(\mathbb{Z})\tau\mapsto\frac{1}{1728}j(\tau)}}{{\simeq}}\mathbb{P}^{1}(\mathbb{C})$ by using the intermediate modular curve $X_{0}(50)$ so that the quotient map factors through it: